OpenStudy (anonymous):

A rectangular table is two more times as long as it is wide. If the area is 72ft^2, find the length and the width of the table. What is the width? what is the length? medals will be given

OpenStudy (anonymous):

w*2w=72 2w^2=72 w^2=72/2 w=sqrt(36) w=6 L=2w L=12 wide=6 length=12

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

you're welcome.